20:00 Cloud overhead dropped rain drops. Otherwise, horizon clearing up.

20:40 we are open! Not photometric.

20:55 LBC archive stuck. No data being put in /Repository.

21:30 green light to reboot LBC archive. LBC archive back on line, but keeps the data from us. Seems to be a mount problem.

21:45 Peter mounted the nfs Repository directory manually on the archive machine. Problem fixed.

21:55 good seeing 1.2" (Uband) but Clouds.

22:05 Lockman U (1.1"-1.2") becoming photometric.

00:15 standard

00:30 GRB seeing 0.9"(R)

00:55 Cluster A2124_U1. seeing 1" (Uband)

01:55 seeing 1" (Uband) A2124_U1 + A2124_B

02:45 NGC 6946 2 OBs in U.

03:45 flats in twillight

a very good night with good stable seeing.

-- NicolasBouche - 19 May 2007
Topic revision: r1 - 19 May 2007, NicolasBouche
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback